OpenStudy (anonymous):

1. choose the equation of the line passing through the point (-1,-2) and perpendicular to y=1/4x+2. a. y=-4x+2 b. y=-4x-6 c. y=4x-6 d. y=4x 2. which of the following is an example of negative correlation? a. time worked vs. money earned b. age vs. commute time to school c. house address vs. math test score d. time eating vs. food left 3. write the equation of the line in slope-intercept form with a slope of -5 and a y-intercept of -1? please i need help on the 3 questions all help is greatly appreciated. thanks in advance.

OpenStudy (anonymous):

1. (y-(-2))/(x-(-1)) = -4 (if 2 lines are perpendicular, the product of their slopes is -1) (y+2)/(x+1) = -4 y+2 = -4(x+1) y+2 = -4x-4 y = -4x-6 (option b) 2. a - the more time worked, the more money earned, so this is a positive correlation b- there is no relationship between age and commute time to school, so no correlation c - there is no relationship here either, so no correlation d - the longer you eat, the less food you have left, so this is a negative correlation - as time increases, amount of food decreases. 3. y = mx + b where m = slope and b = the y intercept y=-5x-1

OpenStudy (anonymous):

thank you very much for all your help